Integration Commissioning meeting minutes - DarkLight-EXP/wiki GitHub Wiki
August 19, 2025
GEM stuff
- “GEM formerly known as Jasmine” is cured, sitting at 3800 on 80% Ar
- No monitoring tool available to be certain of discharges, but it is holding voltage
- Would trip off if exceeds trip threshold, so we’d know
- At TRIUMF: some discharges on LG1 that calmed down. Right sector about halfway now
- LG0 was having 1% higher current this morning, and not smooth. Turned off to let it cool off, then turned it back on
- Really weird looking current plots
- 1% off at any voltage. So maybe the power supply?
- TODO: go swap the HV cables at the power supply
- Maybe later: get new better cables?
Connections
- Did Laura order set of 10 adaptors?
EPICS variables:
- Just have everything. Why ever not?
Readout attempts:
- Should be able to, but we had some readout issues?
- Jan: cannot run cosmics with the trigger right now
- TODO: set up for A trigger
- Cannot run time synchronised without further updates to FPGA?
- But if we have low cosmic ray rates could be OK
- If GEMs look fixed after cable swap we could do this and go ahead with cosmic data taking
TODO for gems: Goal: figure out why current too high for LG0
- Write down what voltages we are at right now
- Top: 3800
- Bottom (LG0): 3000
- Ramp down GEM HV on both LG0, LG1 to zero (confirmation it actually turned off? it seems to depend what order I push the buttons in)
- Turn off power supply - just manually at rack? Anything else need to be done first?
- Unplug and swap cables
- Turn on power supply
- Ramp GEM HV back to current values. Go through slow intermediate steps?
August 12, 2025
Trigger/GEM communication
- DAQHisto working again
High rates in trigger yesterday: Stephi was cycling through targets and talking to Controls about it
- Was sending beam down there? Must be ..
- Update: no, no beam. Beam went only in the low energy area
- Yanyun was messing with EACA, Stephi not sure doing exactly what
Kate to follow up with Stephi today:
- Camera?
- Talk to Victor? Stephi to contact him
- Target ladder controls?
- Stephi wants different labels
- Limit switches not working/not integrated with rest of system
- Getting variables to Jan
- To do today
Spectrometer settings/turn-on
- 192 amps is 30 MeV
- 64 A is for 10 Mev
- Cannot control ramp rate on power supplies as delivered - set a current and they jump there immediately
- Jan has programmed slow control to ramp nicely at 2 A/s. Ernie would prefer 1 A/s but also suggests we ramp to full current (225A) and then ramp down to desired current. Do that every time we change the setting.
- TODO check with Jan if he implemented this
Commissioning spectrometers
- Cabling as only way to switch polarity?
- Electrical safety guys added switch so we can swap positron polarity without uncabling
- Made a stink that only the electrical group can touch these cables
- TODO Kate: go check if it’s there
- Konstantin has had trouble with switches for this kind of thing before, they burn out if there’s any oxidation or whatever
Kate to confirm what’s the situation
- If no switch but same cabling, polarity should match -> this is the case
- If switch, turn on to check -> nope Ethan looking at e-log entry and finding that fixed Hall probes don’t have a sign Our power supply does not have polarity/sign? Ask Ethan/Jan to take a look. Really want to confirm Should measure with a magnet or hall probe
August 5, 2025
GEM tests with visiting experts
Cathode is what’s sparking Tested resistances across individual planes and they were fine Tested cathode layer and that was the cause of the sparking
- Michael: cathode against ground with negative voltage? Yes.
- Did you test to next neighbour? Laura: didn’t test that. Cathode to ground was tripping.
- Jan: other foils disconnected or grounded during that time? Laura: disconnected.
- Michael: did you hear anything or see anything? Laura: can’t see anything, it’s internal. Sounded same as ever.
- At what voltage did it happen? Around 2500, same as before.
Physical GEM installation etc
- Long poly tubes could be outgassing?
- Michael to do some more investigations in the lab
Criteria for turning GEMs back on
- Dust: flush, done. Safe to turn on as long as don’t go over 3500
- Fan vibration: keep off for now
- Grounding: done
- Humidity and leaks:
- Turning them on: are we ready?
- Michael says try one. Don’t do more until we see how that goes.
Mitigation and testing
- Turn on to 3000 or 3500
Procedure: 2000 right away, then 2500, then 3000, then 3100 etc up to 4k. Leave for an hour or two at each.
- Could do just one side at a time.
- Left side top a couple days, then both left a couple days, then top right a couple days, etc
TODOs (physical)
- [TRIUMF] Put GEMs in the mail
- [Hampton] Investigate sparking GEMs on return
- [TRIUMF] Turn on LG1 using gradual steps of voltage, leaving it for a while on each to see what happens. Not dependent on the above.
- [Hampton] Email Rui
- [TRIUMF] contemplate replacing soft poly
DAQ discussions
- Can read out right side at 1.8 kHz from 5 kHz pulser. So at 100% dead time it’s around 2 kHz but if we push harder we get more readout errors.
- 2 components to readout errors: 1) MPD not so happy about long trigger signal. Need to reset the latch to end it, maybe this is a problem. So close that the MPD is confused. Should implement that trigger signal is a fixed length, so there’s some off time before trigger is re-enabled again.
July 29, 2025
GEMs - physical stuff
Step 1: flush line
Step 2: take GEM out, replace
- Test under HV before we put it down
Step 3: send GEMs back. Hampton to learn Rui method, and if it works, teach it to us
- Check GEM window conductivity
Step 4: reassemble.
- Leave the triggers higher
- Keep GEMs together? - yes
Shipping: was FedEx international on the way here, do something similar on the way back? Doug H: need to close the gas inlets and on the outlet apply a filter. Want to avoid pressure changes that suck material into the GEMs or pressure changes or whatnot
- Michael super did not do that on the way here, with either set of GEMs
GEMs - DAQ
Jan fixed half of the readout problems with the firmware, so data loss is fixed. Uncovered a different error that he thinks he can fix. But we have a different problem: while the MPD receives events, it doesn’t write it out (?) Doesn’t receive it back from APV. Only happens in long cables. Looks like the APV doesn’t receive the signal.
- Raise output voltage? No.
- Put a repeater in the line? No
- Custom cables? Slow
- Adapter to other cable type? Maybe but inconvenient Can we just throw away the bad events in the DAQ for now?
- Jan: definitely not long term solution. Maybe OK for commissioning Hypothesis the issue might be the 90 degree connectors or short cables
- Tested: Doesn’t look like short 90 degree cables are causing huge problem.
Take a scope down, put on the trigger, and look at it?
- Which pins?
July 22, 2025
Stephi: need EPICS info Problem with readout: bug in the firmware MPD reads out APVs, then there’s an “event builder” that looks at frame data, finds APV data, puts into memory. Goes first into its local RAM, then info short FIFO and read out. Happening in MPD itself - crate computer reads out from FIFO. If FIFO is full it loses some data. But it’s really small and fills with one event. Meant that big memory in between saves this so you can read out at least a full FIFO every time, but right now we are losing info in the middle of the read. Event header and trailer, but not seeing the trailer. Just going to header of next event. What about MUSE? Events are smaller. Newer version at JLab packs events more efficiently but Jan’s attempt to try it didn’t work well. To try again we need to be able to flash the firmware. Can do either via VME or using some other thing, but nervous to do without a (“JTech blaster”?) to recover it. Device to download the firmware. Checking with KO if he has one Jan flying out tomorrow. Other news: have firmware source code and it compiles, so there’s that. Maybe make the FIFO larger?
GEMs Jan worried it’s not dust because the top two were the ones that started first. Should certainly blow them out the other way Start with the one in the lab Will take a couple of PSI. Flow meter or gas system that Doug sent with first 2 GEMs has high pressure regulator and that feeds into safety valve and then low pressure regulator that outputs 1 or 2 PSI. GEM can take 1 to 2 to 3 PSI no problem. Take off bubbler for flushing so we can get more air through it. Get a 10 micron filter to put on the end of the line in future If there was dust on the line, why only 2 GEMs showing trouble? Why not the HV reverse bias? In principle, if you reverse bias so you charge the readout layer, it pushes any filaments to have the opposite polarity. One of reasons for sparking is there’s a little copper filament that didn’t get cleaned off and it can burn off, but if not it’ll just sort of get into position and stay there. Rui’s high voltage shock system sometimes works and sometimes burns things out. What Doug and Ethan did: disconnect high voltage distribution and then you can apply voltage directly between 2 gem layers, or gem to readout board, or whatever. And do it in both directions. Rui didn’t like this, thought it should only be done in proper direction. Rui is real expert and says he doesn’t like this
Taking them apart: problem is tensioning when putting them back together. Need someone who knows to show us. That’s only Michael. Can we prioritize getting Michael to train us when he’s here? Does Michael have a cleanroom now? If not, makes more sense to get him up here to fix them rather than sending them back. He has device Rui uses to apply shock.
Alternate (non-dust) hypotheses:
- Fans? Left one started sparking soon after we installed the fans. The only thing different between lower and upper GEMs is presence of fans. More vibration on upper than lower, and some air on the tail end of the window. If the window pumps up and down, and if we have large oscillations on there, it provokes sparks.
- But then why would it continue when we take it off? Could be sparks are so bad that something broke and now it’s continuing to spark on the table.
- Could build something with a bottom on it. Can’t hurt.
- Diagnostic: turn off fans and see if problem goes away.
- Two grounded foils oscillating near each other?
- If you press on outer foil it increases pressure on GEM foil, which makes the GEM foils near each other. Dynamic change on volume above GEMs could do this. But it’s the readout plane that’s oscillating so that’s probably not going to do anything.
- Humidity? More humid than usual last week. But no problems observed in GEM lab. Would only make sense if sparking was external. And HV probably not high enough. But HV very close together …
- Grounding? HV was applied before grounding was done. HV ground is upstairs (?)
- Jan: cable of high voltage is not grounded. Power supply upstairs is not strongly grounding.
- Only way this could account for sparking is if sparking was from one GEM to another. Wouldn’t account for sparking continuing when moving one to table, but two GEMs might not be sparking for the same reasons.
- Gas dirty?
- Leak in the line letting some air in? We didn’t pressure test it.
- Test 1: absolutely have to take over to GEM lab and see if it still sparks if it’s on the GEM lab gas system, humidity, etc. Let’s see if it’s something in the hall that makes it spark.
- Could be a bad connection in one of the Ts?
- First step: move back to GEM lab. Go with gas system, known tubing and location, blow some dry nitrogen on the voltage divider (?), test if sparking continues.
- Go ahead and do the gas reverse, with Doug’s pressure limit of a few PSI
Jan: just need more eyes on the GEMs. Move other misbehaving one. As long as APV cards are plugged in, powered, connected to ground, then we can check if it’s sparking. Need: LV, (not HV?), might be enough to just plug in?
Flush in situ, test now that grounding is connected? Or move?
June 17, 2025
Lead cutting?
- Shears proved challenging - did not cut metal
- Check Rick’s videos?
- Got a good chunk of the way there by hand. GEM front shielding done and halfway done the target chamber shielding
- Punch-out set for the bolt holes
Ethan status
- Ethan and students removed HDMI cables and cleaned up near the rack so we have all the HDMI cables in a corner of the counting room
- Decabled the spare HDMI that was in the cable
- Ran the magnet interlock control, they just need to hook it up
- Other GEM frame: we are missing like 2 boxes of bolts. Tons of nuts, just missing the bolt boxes
- Run new HDMI cables today
Yes send back Michael’s power supply
Transducer’s been installed. Not fully cabled up yet. They are there but not connected - need to give measurements to controls so they can get the target ladder locations right
Cables connecting: just need to talk to electrical group
One steerer to do today, Hubert needs to hook up power
Vacuum: still pumping down. Was at 10^-5 on the weekend and wanted to see if they could do better, but haven’t got the ion pumps going yet. This is just with the remote pump.
Mechanical services: Paul’s missing fittings? Can they run without that? Probably
Jan’s MIDAS not running by default Dl-server-2 Username daq Password DL17 or DL2025, can’t remember
TDC cable problems:
- Not seeing calibration pulses
- Some signal attenuation but we should still see them
- Mike: calibration pulse comes in on the i2c cable. Peter and KO were both puzzled about this
- Damage to distribution board? Some active component burned out.
June 10, 2025
Vacuum stuff probably wait til tomorrow Craning: do the hard block today and second one tomorrow
Outstanding stuff:
- Water & air: ideally tomorrow to finish
- Controls:
- Hubert will do modifications once target ladder is in
- POT arrived yesterday so we have that. Do that at some point today
- Other cables in, need to hook up
- EPICS updated; Laura and Kate getting access
- Do we also need access to HLA? Sure. Stephi to set us up
- Doug: what surveying is Tony doing?
- Wants position of collimator, magnets, and exit flanges
- Laura: did that order of things in aligning first arm.
- Tony will give us the as-builts when everything is done. Doug wants as soon as possible; needs to know where the exit flange is so we can tell if it’s parallel to focal plane, high or low, or what
- Doug basically just wants to know if it’s above or below the focal plane.
- Laura checked collimators are correctly aligned
- Doug to Jim and Joe: what do collimator survey buttons line up with? These are what Tony will be able to actually survey
- Joe checking
- Buttons almost aligned with front face, but a 1/4” offset with the buttons (buttons are closer to target)
- Doug assuming Tony measuring centres of buttons? Laura: one button and flange faces.
- We need to clamp one of the parts of this support structure to hold it in place??
- Doug to Jan: you’ve been programming EPICS for magnet. Is there an algorithm?
- Jan: ramps rather slow. Interlock will switch off fast.
- Doug: let’s say want to ramp to 200A. Thinks should ramp to 220, leave it, go to 190, etc
- Jan: planning to write something with a feedback loop based on Hall probe measurements
- Have to control by Hall probe anyway since the readback on the current is pretty low precision (certainly compared to ability to set it)
- Should we add extra lead in front of Hall probes to protect them from radiation? Add a shot bag on top of the chamber in the right spot?
- Need to run the Hall probe cables before we can test them downstairs
- Add a lead hat to the target chamber?
- Has Joe shipped APV supports?
- Gave box to Ann on Monday afternoon, should go out today
Best way to cut 1/4” lead?
- Doug H: shears?
- Jim would use a wood blade with water cooling
- For the holes, a press with a pipe as cookie cutter? Jim: dunno, could try it. Would dull your cutter. Jan: make a one-time cutter from aluminum pipe? Trashing it is fine
- Doug: don’t you have 1/8th? Much easier than 1/4 to shear
TODO Kate:
- Go down and check if we have any bigger pieces that are 1”.
- Ideally we get three diameters and let’s go see what they have
June 3, 2025
Laura to Jim and Joe: why does fiducial on arm bellows seem to be different from drawing? Did they put the collimator in?
- Jim did not put collimator in bellows.
- Joe says sounds like collimator isn’t aligned properly. Checking model.
- Conclusion: need to take out the collimator and rotate it. Laura will talk to the TRIUMF guys
Once all beam pipe components connected, will reinstall target ladder and start to pump down.
Target ladder:
- New transducer to arrive Thursday
- Can transducer be added with the ladder in situ or does it have to be done somewhere on a table?
- Requires that a mounting bracket be made. Ethan can make it, says Doug. Needs screw sizes and locations of hole
- Can certainly mount the transducer in situ. But will it connect properly?
Needs to know location of top two screws, screw size, and distance to hole in the middle 
Ernie says we can tell what thread the old one has by just looking at it online and we should be able to figure this out. Then we can just make a new part to adapt. Tony may have made measurements relative to this adaptor part? Laura: measured relative to three-pronged piece. It won’t change so we are all good
This gives us reasonable confidence we can do this after installing target ladder.
Kate to come over to counting room after this and to bring snacks
May 27, 2025
Spectrometer field
- Max current 225A
- Makes 0.4 T field
- Jan working on slow control
TRIUMF status update
- Target chamber and 2nd dipole went down yesterday, being aligned today
- Rest of beam pipe is ready to go, minus collimator section
- Lucas wants some more practice rounds before doing the real deal
- This is the bottleneck now
- Brian working on the GEM supports
- Electrical services is working on power cable for Jan.
- Are covers for the back of the dipole power supplies still in the boxes? Yes
- Laura will give them to electrical services
- 480 phase. Outlets in the rack already
- Cables going to the wall. Installing yesterday or the day before
- HV cables for GEMs probably need to drop sooner but rest of cables should be good to go to the wall
- Positions of cable trays? Mike and Doug P have discussed something but no one here is quite sure of the arrangement
- TODO Mike: get Tony the drawings for the holes on the back of the dipole magnet. Mike will do
Jim on seismic safety:
- Scattering chamber support is in two parts, top and bottom, to allow for adjustment. Two parts aren’t attached to each other. So once it’s in place, connect them. Jim was going to use stainless steel tie rods, thinks that would work. Or a steel packing strap. Stores has stuff we need.
- Four ties, top horizontal bits to those in the middle
Doug H: did you re-word [document] in way that you wanted it worded?
- Doug says purpose of the document was to get documentation for TRIUMF. Thinks Josh should sign it.
- Jim and Doug think liability is implied in some of the sections.
- Laura: can we re-word it? Doug: [silence]
- Jim sent a re-worded version to Doug; has been told he cannot sign it, it has to be Doug and Richard Doug TODO: read Jim's edited version and sign it.
Ethan: new transducer, find out where it is
- Doug recommends switching over to a resistive transducer if they can't find it
- Question about time frame
- Newark thought they had some in inventory but they can't seem to find it so we'll see
- "I'll see what happens"
- Jim will look around and see what he has
- Doug has found resistive POTs for $200 or $300, but Jim says if he has something he could send it out overnight
- Not a 1:1 replacement, would require a new mounting bracket and some changes to connections and power supply. Doug could do this but isn't going to be here. Four wire connection to a 3 wire connection, so have to rewire the inside of the power box. Fairly straightforward, just have to make the ground common between the 24V that's the excitation of the resistor and the potentiometer that has the ground
- Solder and re-solder a couple of wires, one for potentiometer and one for remote control
- Brian could do the new mounting bracket
- If we can get the POT to Brian before he leaves, then he could probably make the mounting bracket and figure out how to connect it with the existing target frame
- Might be faster than the replacement if the supplier really can't find it.
- If they can find it, that's top priority and our #1 choice. Only thing that needs to be done is Hubert needs to add the connector. But if they can't, then it's good to have backup. Jim will check right away.
- Laura will hold off on target ladder as long as she can, but it'll be end of this week or early next week, so the amount of time we have is very little
Laura: Targets and stand-offs on the way?
- Doug: should have arrived last week
- Laura only got the screw package
- Ordered from McMaster-Carr. No shipping number of expected arrival date. Jim will look into it if Doug sends it to him
- Laura will check at Stores.
- Laura prefers to put only one Carbon since there is only one spare, and if we break it we don't want to break two at once
May 20, 2025
Last week: getting the tables aligned
- seems to be going ok Target chamber being leak tested today
GEMs: disappeared into UPS
- Show as “investigation opened”
- Did we test power supplies for dipoles? No, Shaquille used his own
Ernie wants details of cables to be installed: type and length. Laura has it
- Don’t think we can reach the topmost target position. Need to swap carbon & tantalum target when we switch beam lines.
- Could we make the ladder longer? That’d be quite a bit of machining.
- Jan: what is the limitation? Doug: maximum throw of the screw doesn’t get it down far enough.
- Various ideas passed around: put in a spacer to get targets to sit lower? Modify? What about the bellows?
- Laura: we should leave this until July to fix because we want to leak test it today. Let’s design something better.
- Do we need the carbon target in July? Yes best to still have access to it.
- Can’t we figure this out from the drawings? No, the drawings are wrong.
- Measure beam to flange in the old chamber, compare to drawings of new chamber? Old chamber blanked off but could measure on the outside
Dough H: Transducer on target ladder not working. Need to replace it. Discussion re the transducer. what are we going to do about it?
- $1300 or 1400 for an exact replacement, or something cheaper that’s approximately the same thing. Ernie: don’t need same resolution for it as for original application
- Could new electronics be a step counter?
- Ernie: it’s more annoying, you have to reset at the end, account for screw, etc. Better to replace LVDT.
- MIT decision which one you want to buy, but just go ahead and buy one and get it done.
20 degree magnet aligned. Next up is scattering chamber.
- Will leak check happen on the roof? Yes in cleanroom. Then we’ll bring it down afterwards.
- Jim says get whole unit pumping together as soon as possible: going to take forever to pump the PEEK out.
- Doug H: don’t want to put on the 36 degree yet cause it gets in the way. One thing at a time.
TODOs Kate:
- Pester HR for Jan’s A-number
TODO Doug H:
- Fix transducer for motor. Up to you how.
- Decide an approach to figure out how to modify target ladder for the top target position; make it happen
TODO Laura:
- Send cable info to Ernie
- Get power supplies plugged in so we can test them
- Talk to stores re customs brokerage for GEM power supplies (and GEMs?)
May 13, 2025
- TRIUMF
- Beam pipe is fully assembled
- Removing old target ladder and replacing the 10 um Ta to 1 um C
- Needs to be aligned, will start shortly
- Going to move the spectrometer arm down to the e-hall
- Package received at stores, along with lots of 80-20
- GEMs
- Supposed to be shipped on Thursday
- Also need to sit on nitrogen for several days
- Manju arrives today
May 6, 2025
- Laura’s update, not attending given to Ethan as text
- Hatch opened yesterday, dump shielding is being removed today so we can access the eLinac collimator location
- Vacuum leak checking is underway for the arms, assembly yesterday, will get to leak checking today
- Various issues with the integration were found and are being resolved. This included the eLinac collimator not fitting in the dump shielding with the current design for that section of beamline, the spool pieces downstream of the DL DS bellows not existing, and an issue with the configuration of the vacuum diagnostics box.
- For the collimator: Tony is redesigning the beamline section, will have to get some adapters machined and the piece welded together. Collimator itself should be in hand by end of the week.
- Beam spool pieces: resolved. They will be welded soon
- Vacuum issue: in talks with vacuum group, have workable solution just need to update drawings
- Ethan
- Roger at Rad Safety reached out to me about a CNSC question regarding dose on beam components and experimental equipment.
- We have thought about dose on detectors, hence our shielding. Have NOT thought about dose on beam line equipment
- Roger cc’ed Brad and they will get back to Ethan about what they need
- Ethan is concerned that we are still technically not approved by CNSC, can we run? Suggested to email Kate and ask to get a straight answer
- Tony
- Shows new spool pieces due to collimator being inserted in dump shielding
- Some discussion about getting an ion pump near the dump. Ernie mention that NEG pumps don’t pump noble gasses well, so an ion pump might be helpful
April 29, 2025
- APV Supports
- Doug has redesigned and sent drawings to Jim
- Jim will find out if MIT can make it today/tomorrow
- Rad Safety
- Steffie wants hard numbers on residual dose near target after beam off and target activation
- Ethan will provide those numbers. Everything will be safe
- EHall
- Is now open, can begin moving things down into the hall once cleaned and tested
- Keep in mind it is dusty
- Can’t bring down final dipole until it is mapped, others are good to go
- Bates
- Will bring some of the 14 inch conflats for the vacuum chamber
- Tools will be provided by TRIUMF.
- Bring safety boots!
April 22, 2025
- Laura has collimator in COMSOL
- Making progress in cleaning DL chamber/spectrometer/etc.
- Sending off work permit requests in the next day or two
- Still on schedule for disassembly in first week of May
- Shaquielle should be back on Thursday to map the next magnet
March 25, 2025
- Laura received windows yesterday, also met with electrical services last week to get the ball rolling on that
- Jim: analysis and vacuum items? Vacuum cleaning should be fine, working on seeing about resources.
- Timeline: don’t see any problem with beginning of May. PMOG QRPP only possible bottleneck.
- Doug H: how long is disassembly? Under a week.
- Target ladder: microswitches put in? Fair amount of work, should figure something out for this in advance. Include switches to the controls group.
- Hot filament cable: need a long cable for hot filament gauge. Talk to Nagma and Dimo to see if they have a solution. Doug H sourced it from the company and it was $3000, will try to find another solution
- Stephi: planning to talk to Holger to get cables for collimator, and working on controls for collimator. Laura will follow up with Yetvart about collimator welding.
- Doug H: Ethan will be here first week of April. Do we need people before May?
- Magnet mapping: don’t have the rated plug for the MIT power supply. Existing power supply will work that’s already on the rack.
March 18, 2025
- Tony: informal conversation with vacuum group, they seemed optimistic about getting the cleaning done in time. Won’t happen in the next two weeks, but should happen after that.
- Jim: How does moving it down into the hall work? Do the cleaning, cover ports with UHV foil, then move down into the hall. Alignment: we can help out. The more the merrier in terms of hands-on help.
- Stephi: should we have the hall cleaned in advance? Tony: maybe, not sure whether we have/need a tent and stuff.
- Doug P: might be a big tent from the RFQ repair. Quasi clean room, only keeping dust out.
- Shaquielle is pretty overloaded so probably can’t do much with the magnet mapping right now.
- Doug H is at TRIUMF.
- Stephi: had 3 days of continuous beam delivery last week at 30 MeV, a few trips but 98% up time.
- Ethan: magnetic field, had done shift by eye but can now do a fit since we have resolved the jump in the field. After doing the fit the agreement is generally much better, in the pole tip have to 1% agreement. In the fringe fields still see some differences. Fit thinks we are off by 3 mm in the y direction. Think the probe may be at a slight angle going into the dipoles. The scale is also a bit off. The measured is at 200 A and the calc is at 192. Around 6% difference, should only be about 4%. Tony will check with Shaquielle about what the exact procedure was for when he checked the current. Ramp up may be part of the issue, will check with Shaquielle.
- Can just do one measurement between the pole tips with the MIT hall probe and power supplies. If it’s still off by the same amount we know this is not the issue.
March 11, 2025
- Laura updates: vacuum parts from the crate are being inventoried and then passed off to vacuum group to clean, productive meeting with electrical services, should be able to do what we need.
- Jim: updated talk uploaded and window test video
- Josh: ANCOR well-established manufacturer, may need some weld information but not sure this is really that important. We’ll see what Josh actually wants before pursuring this further
- Discussion of upcoming QRPP.
- Laura will discuss with Ritu once she gets back from vacation
- eLinac: running at 30 MeV, everything looks really good. 3 kW power.
- Working towards running longer term. License enforcement system, needed to run with the collimator.
- Discussion about commissioning plans. Decision to run as expected as we can do target out runs with the 30 MeV beam and the low energy configuration
March 4, 2025
- Crate is in the process of being unpacked, Laura is inventorying everything for vacuum group that needs to be cleaned. Tony is checking with vacuums to see what their load is and whether we can get a clean room.
- Collimators for arms -> Jim has concerns because it’s tungsten and maybe shouldn’t go in the ultrasonic cleaner.
- Jim also would like to talk to whoever is taking out the PEEK.
- Laura has various updates: haven’t heard back about power cables, mu metal has been located from Thomas, will follow up with Shaquielle about getting it cut to size. Have sourced all parts for the 2nd PMQ spool piece so that will be good to send to UBC to be welded once the PMQs have been mapped.
February 25, 2025
- Ethernet cables: may need some for the trigger, Mike passed to KO
- Stephi: had a chat with Mike from controls. For this QRPP cycle we will not get any help from electrical services. Will have to fight for this next cycle.
- Questions about the B field: more than one file will need to be stitched together to get full map with the fringe field. A measurement takes around 26 hours. Need measurements on the inlet side or the GEM side? The GEM side is more interesting. Doug H still thinks it isn’t useful to do the full field map for tracking. The step: HAICU’s work permit has expired, so can’t test if that was the size of the step. Vertical step is a bit weird if it is another magnet, power supply resolution may not be sufficient? COMSOL simulation? It looks fine, but doesn’t map easily. Ethan has asked Holger to calculate with a 1mm grid instead of voxels. If Shaquielle has time, redo measurement and see if it was indeed the HAICU magnet causing the jump.
- Technical package: seismic analysis has been running for the last week. Vacuum results can be up tomorrow. Laura will set up a meeting with Jim, Joe and Josh.
February 20, 2025
- Tony: Additional PMQs received, Shaquielle is continuing magnet measurements and working on figuring out what’s going on with the interference and the power supply resolution. The frame designs did not include grouting, might be a bit tight to accommodate for uneven floor.
- Doug P: looked at floor data, added enough room for feet and stuff. If floor is too high we can trim leg, if too low we can add grout.
- Stephi: during the installation, we have to have an account to pay for stuff. Can we use the regular grant account or do we need something separate? Can just use the PRJ account.
- Doug H: crate left this morning from MIT. Should arrive next week.
- Doug P: can put crate in the proton hall extension. May have to put some stuff on the ehall roof.
- Doug H: may come out when he gets back from Germany (around the 15th to the 23rdish).
- Should we get timelapse camera for installation. Laura can ask Jesse, Doug P also has one from ARIEL target we could perhaps use for a while.
- Laura: we really need the technical package for gate review approval and engineering approval. This is becoming a bottleneck
- Magnetic field mapping: 3D would be quite long as would have to sweep twice. If we really wanted this, would require someone at TRIUMF to learn how to use the machine to run these.
- Probably the HAICU magnet causing the interference, still need to confirm.
- Really need to confirm what we actually want. Still want a good fringe field measurement. Need to decide on exact step size and how many planes we want. Might be able to get both fringe fields off one map, Doug P will check.
- Got the field map from Holger from Comsol. Format is a bit weird, will get Holger to read it out in the same coordinate system as the current map.
- Laura: Do we have a spare piece of beam pipe for the new PMQs? Doug P will check, otherwise we will need to order.
February 6, 2025
- Went through todo list, some notes and updates added there
- eLinac: cold box tripped so warmed up. Now cooling back down, nothing broken
- Had 18 h continuous beam before having a small trip (yay!).
- Hoping for three days of continuous beam next week with 27 MeV
- Then 30 MeV for three days post that
January 30, 2025
- Joe is back, was helping out with getting the crates packed up.
- eLinac: swapped amplifier last week, the old one was dead. In the process of getting a new one ready. Working on running with the spare. Just broke due to old age. Something was going on with the electronics, were getting frequencies that were not expecting. Beam current is relatively stable, trying to do some longer runs before ramping up the energy. Stephi is pleased :)
- All CNSC docs are on Roxana’s desk. She’s working on getting them out.
- Filled in remaining to-do list items
- Laura: checked up with SABR, second set of PMQs should be shipped next week. Collimator is in progress still, cooling tube needs to be bent into shape. Dropped off at the machine shop and then will bring it back to ATG.
- Power supply cable follow up: Doug P chatted with Marco and he has used most of what he had, will have to buy more through electrical services. Tim, Stephi, and Laura will chase down electrical services.
January 23, 2025
- Stephi had meeting with vacuum group, re integrating into EPICS. And met with controls. Needs a few things. Vacuum group, do we have plans for diagonostics around the chamber. Jim: we are using eLinac gauges. Have a free space where the IR camera. Jim: doesn’t think we need gauge on the chamber.
- Combo ion and neg pump: do we need two tables and readouts? Jim will get info on how to activate integrated neg. Jim is checking if we have an extra neg module for extra. Pump is underneath.
- Organizing a meeting between Jim and vacuum guys so they can talk directly
- Mesh between chamber and ion pump? Jim will check but he is pretty sure it doesn’t but that it is not needed.
- Doug P: usually don’t do on ion pump. Just on turbo.
- One’s a convectron gauge and one’s a hot filament.
- BPM body is now in front of the chamber
- Who is installing foils? Either Doug P or another group that is used to foil installation. The thin one will be fussy.
- Target ladder: has BeO on bottom.
- Camera is not completely in focus anymore, Stephi was hoping we could maybe use it for beam size measurement, but may not be possible
- Doug P can maybe add the visual alignment thing to the target ladder frame at the same time as the target is changed.
- Crosshairs on the screen would screw up beam size measurements.
- On ladder instead of frame? Probably better.
- Make sure we check what position the targets are at. Doug P has the order somewhere
- Signals from spec magnets and Hall probe will be on DL system. No cross talk with eLinac? Jan: why? It is in EPICS. Stephi: limitation from controls group. Controls Mike doesn’t want cross talk.
- Stephi will setup meeting with Mike (controls), and others to talk about setting this up.
- Flowmeter for water for magnets? Doug P: we will need that. We are providing it as far as Doug knows. Will have to go through Proteus. We would need to be responsible for the flow meter on the controls side. Will talk about in meeting.
January 9, 2025
- Joe & Doug discussion
- Asked about numbers for installation drawings, Doug is following up with alignment guys that what we have is fine
- Also commented on PEEK and how it can work with cleaning: Doug properly forewarned
- Each dipole chamber has two “PEEK pushers” - rod welded to conflat that pushes PEEK into position. Needs aluminum conflat gaskets
- On the magnetic field mapping
- Doug to bug Shaquille next week, he’s busy with ISIS
December 19, 2024
- Installation or assembly drawings: have to create these
- What are they looking for? X Y and Z views from given point? Is the crossover point good or do they need the other point?
- Questions to Doug P:
- Typically for installation drawings we look at all of fiducial points back to cyclotron center and ask for their locations relative to that
- Plan view for fiducial points and we do x, y, and z locations of each fiducial point back to cyclotron center?
- Doug:yes. Because there’s a lot of ambiguity when we pull those points out of a model
- Asked for signed drawings for installation. Before, people have said “oh it’s 20 inches from the center of the device” but nobody defines what the center of the device is
- Joe: if you have the fiducials from the 2 dipoles and the ones for the scattering chamber we have all of them
- Doug wants to talk to our alignment guy Mike and see if we can set up the crossover point as an origin for this
- Joe: if we get the origin we could also scale from that
- Doug: but we do know where the crossover point is with regard to the cyclotron center, so that’s good enough
- So installation drawings referencing the crossover point is fine
- Summary: that’s like 12 points, not so bad …… if Doug can confirm this with Mike the Alignment Guy, Jim and Joe can take a look at it and see how much work it is and go from there
- Mike confirms that crossover point is fine!
- Joe: CAD model would be accurate past that
- Time estimate to do this: on the order of a day. Doug H: manageable!
- Doug P will join the chat this afternoon
- Doug explains to the MIT guys why we are so careful about this!
Who’s going to make the frame for supporting the GEM shielding?
- It has welds, so probably not SciTech
- UBC shop?
- Doug H: rather than weld it, why not put a plate with screws at the right points?
- Some discussion
- Jim thinks it could be replaced with plates & bolts
- Doug: typically these extrusions are anodised, and then you have to take all the anodising off to weld it
- Jim: that’s not bad actually
TRIUMF to make cut shielding
- Make sure TODO list item is there for getting the stands made
- Negotiate between SciTech vs UBC shop to get it done
- Doug will do the lead cutting and shaping and painting - has to do a ton of it shortly anyway
MOU
- MIT happy to loan stuff and insure it while it is in transit and here
- Can’t insure against damage to TRIUMF if something bursts into flames
- Deal with it at the collaboration meeting
December 12, 2024
Is drawing up to date?
- Yes Can we add steerer?
- Okay Project management
- We’ll switch to Microsoft Project
- Doug and Stephi to import spreadsheet into there for a starting point Cleaning
- Could do with PEEK in there if we have to, which it looks like we will have to
- Could do a low bake depending on melting point of PEEK to get the moisture out without freezing it
- Jim really shouldn’t bother cleaning everything cause it’s all going to be recleaned here anyway Consoling Dimo about the vacuum
- TODO Kate: email Jim and Dimo to suggest they have a call Magnetic field mapping
- Doug and Shaquille taking another look
- TBD!
November 21, 2024
- Joe making a list of bits and pieces. Going to send us spare copper conflat rings, windows, etc
- Doug will assemble list of things they will ship
- Including identifying which long cables he has. Thinks it’s for the convectron but not sure. That’s one of vacuum gauges but there’s another one also.
- Other thing that will need long cables is Hall probe.
- Already bought them - done
- Ion gauge: Doug will send it, and should have a controller, but we don’t have long cables for it. Need those
- Power supplies arrived
- Drawings
- TODO Kate: follow up with Stephi re location of camera box
- TODO Kate: follow up with Bill Richert re engineering sign off
- Mike will email Doug P asking if he has the beam pipe.
November 14, 2024
GEM measurement
- 18.2 mm?
- Exit flange: Took down to 12.5 mm and decided to counter-bore screw heads so they wouldn’t puncture the foil
- Found some diagrams in Michael’s tarball that make it look like it should be 17.4, so bigger than 15 but smaller than Laura’s measurement
- What happens if we move the spectrometer? Can we get the focal plane where we need it? Doug will try moving it around a bit and see what the impact is
- Otherwise we’re just going to have to do the best we can with what we’ve built
- Not worried about squishing the bulge from the GEMs - we’re already planning on that.
Doug worried about TRIUMF saying they want to clean everything
- Getting the PEEK out is tricky
- Does TRIUMF have a whole supply of the right size rings etc? Other flanges? If we need spares
- Doug H should email Doug P and ask
Beam pipe PMQ spool piece
- Doug needs to track down some longer beam pipe
- We have everything else now - Kate can get PMQs to Mike
Controls and cables?
- A mess
- Konstantin says talk to Rick about cabling for low voltage power supplies
- Whole bunch of cables in the electrical area that look like we could use them, maybe. So ask Rick
- Need to talk to Doug about cable tray. If we go down Hole C we have a straight path. If we go down Hole A we need to move 2 m perpendicular to the wall and then we can go straight to the spectrometers.
- Go masking tape out the spectrometer locations on the floor?
- Don’t gain a lot by putting in a diagonal cable tray so we can go to the wall and that’ll be OK. At the spectrometer, height of top of trigger scintillator boards is 56” over the floor
- KO and Mike will set up one trigger panel downstairs in the next weeks and power from upstairs and see how it works. Should get short periods
- DO NOT FORGET: cables for the vacuum gauges??
- Ion pump, two gauges (convectron and hot filament)
- What do we need? Doug bought one 50m cable for one of the gauges. The other one he didn’t buy, was hot filament one with plug in to unit that already exists. Thought the statement was TRIUMF had these sorts of things.
- Hall probe cables are done, Laura bought them
- If the shipment for the power supplies hasn’t gone out yet, add the Hall probes to it
October 31, 2024
Power supplies and cables
Flanges making their way over
Sorting out measurement for Doug for field mapping
Joe:
- Got downstream bellows spool piece drawing done
- Going to get it welded at the same time as the chamber
- Back to supports for GEMs and triggers
- Jim finished connection to dump; Joe will update solidworks models on Dropbox and we can get them ECO’d
Ernie
- Sending us the power supplies
- Running at ~240 A
- Hall probe mounts? Not gotten to them yet.
October 24, 2024
Doug H: news from welders?
- Guy is out sick so not clear
- Went down to MIT central shop and spoke to the manager. Found the PEEK and he’s going to start on that in the next few days. Doug told him he wanted it as soon as possible and that we need it before pumping down
- Also need flange for the exit window to clamp the material in place, but he didn’t know anything about that
- Joe to clarify drawings/communication with them and make sure it gets done.
- Spoke with Jim about who they would send to TRIUMF for installation, Chris might be interested. Or Jim if it’s in the new year
- Vacuum components should take about a week to install
- Ethan will probably come when detectors need to go in
- TODO Kate: talk to Thomas about order of operations for controls/electrical
- GEMs locked in place? Threaded rods give us some wiggle room but not enough
Beam pipes here at UBC
- Need to discuss clamps with Doug and see if he is happy or if we need to make modifications
- We need to get the flanges for the pipes and give them to UBC. Need to confirm which needs to be rotating.
- Ask for deadline end of November. Need time for alignment.
October 10, 2024
Discussion of energy measurement issue in integration meeting
- Need to know the beam energy!
- Doug is curious why we can’t get something better from the machine
- Time of flight? Can we take a pickup just downstream of the RF and then at our target or whatever?
- Mike asked a while ago and heard not much enthusiasm
- At 90% duty factor we could probably resolve which bunches we’re picking up How accurately do we need to know this at lowest energies?
- At 10 MeV, +/- 20% is 2 MeV
- Don’t trust magnets to better than 10^-3 in the first place
- Put some toroid pickups at various points along the beam and do a time of flight over 10 m and see what we get
- Do we have capacity to do ns level measurements? Would need sub ns timing for this to work, which they almost certainly don’t have.
- Conclusion: time of flight is out
MIT update
- Joe got the bellows in the mail
- Will be a solution it will just be expensive
- To call Doug P over the holes
October 3, 2024
Lead in back of spectrometer
- Laura sent updated file to Jim; Jim thinks it looks fine
- As long as it’s under 100 pounds it should be fine
Jim update
- Bellows will arrive next Thursday
- Waiting on info from Doug Preddy
Stephi update
- Get rid of BPM between EMQs and replace with isolation valve
- We do have a spool piece we can cut up that has a BPM and bellows in it
- Piece is 20 inches long
- Doug Preddy thinks we can use existing pipes as they are.
- Jim can try to reuse pieces we have. If it works great, if not we cut it up and reuse pieces
- TODO Kate: find soldworks for valve and get it to Jim - done, he has it
- TODO Doug P: send dimensions to Jim for table and location as soon as possible
Ethan
- Might be able to reduce shielding around detectors!
- Steel itself of the magnets offers some non-negligible shielding
- Might need to stick some more shielding in in real life though. Front of spectrometer shielding won’t be as good as it is in the model, and there are some weird pieces in real life around the bottom of the detector shielding at the front.
- We should try to increase shielding at front of detectors for safety. Make some inserts under the brass piece to increase frontal shielding
- However, can probably remove edge shielding. Concern about back scattering from wall into positron detector seems not likely to be an actual problem.
- Jim: work up a preliminary model and after we have some data we see what we have to do
SAR shielding
- Remove detector shielding from SAR modelling and text. State that what we show is a worst case scenario and extra shielding will be added in practice based on commissioning results, but we will not include it here so that SAR shows worst possibilities
- TODO Ethan: rerun without detector shielding for updated doses.
- TODO Ethan: also add new areas to SAR?
- TODO Kate: modify text and images to get rid of shielding around detectors and instead explain that it will be added later and is expected to make a negligible difference, but any difference will be in the direction of decreasing dose
Discussion of frame/support between Mike and Jim
- I think they remember what they wanted, it is not clear to me
September 26, 2024
Clamps
- Jan has done first print design
- Will send to Stephi
- Brass threads in insert
Bellows
- Stephi will send to Jim
- Matches model from Doug
- Joe thinks he has a solution that works using them. He’ll be back on Tuesday and they’ll give it a try
Beamline
- Jim will meet with Doug P next week to check over assembly of the beamline
- Drawings basically done
- What’s missing is Doug P’s information on as-built stands for diagnostics box. Should be incoming.
- Then they will send us the drawings and we can proceed with ECO signoff
PMQs
- Will be ready to order once Stephi has been able to take a day to do the analysis on this and get it done
- Mike wants a placeholder drawing for UBC for the pipe section so he can get in the queue at UBC
- Do we have BPMs? Yes.
- TODO Stephi: follow up with someone to make sure we have electronics for the BPMs
- TODO someone (Laura): go take photo of each end of a PMQ and send to Jan to confirm it looks like drawings
- Jan to send 6 clamps which will need to go onto the pipe along with the PMQs before the flanges are welded
- 0.265 inches to front of first PMQ - can indeed fit a tiny clamp in there
- We should buy ourselves a little 3d printer for making this kind of thing
Support for GEM/trigger scintillator
- Jim: would prefer to talk about it next week
- Mike: asked Philip for a new drawing. Will send it as soon as we have it.
- Jan: would change the L bracket on the side to 80-20 (T channel frame). Have one run along the whole top …
- Jim: will consider it once he sees what Philip has.
- Jan: when we really go upstairs with all the electronics, we have a grounding issue. How do we have a separate ground here between the electronics and the magnet?
- If we were actually in the hall it might be OK because the grounds would be similar enough, but here they aren’t
- Does this allow that?
- GEMs are isolated by design
- What about Al bars on side of trigger scintillator? Only touching plastic
- Fans for cooling
- Absolute total is < 20 pounds and that’s in the noise as far as Jim is concerned. Can support them pretty easily
- Doug: would tend to put them on the supports for the lead shielding.
- Are we worried about them vibrating? Unclear. Michael doesn’t think they add a lot of vibration
- On Wednesday, discussed cutouts where the HDMI adaptors sit
- Cables will have rubber around them at 2 or 3 times the height of the socket, so better remove the yellow part in that area. Not on the GEMs that we have here.
Response about the hut
- Michael currently has six 20-m cables lent from SPS (yay!)
- Took a reference run with standard cables
- Working out reference info and will then swap. Ryan has replaced cables of top GEM and just finished tuning clock phase for that.
- How real is timeline?
- Be as fast as we can
- Mike TODO: poke Konstantin about the same issue for the trigger
- See about 10% attenuation in reference so far; about as expected
- Other piece of information they need is about performance of angled cables
- On the side will have angled cables that run along the side and go to the top corner
- On the short side, can get away with some straight ones.
- Doug agrees
- Lead is supported from the floor. Could add a bar that gives somewhere to support these.
- Joe’s got something - need to talk with him.
- For next week’s meeting then
Project management plan
- Stephi has made a wonderful spreadsheet currently running til end of Jan
- Will send to group managers and people with hats
Doug
- Keeps getting messages from Pauline Della Zola - does he need to be able to access these?
- TRIUMF review process: need to make sure anyone on the distribution list has a computing account that is currently active
Ethan
- Where are we at in the SAR discussion?
- Roger looking into it
August 27, 2024
PMQ mounting discussion
- TODO Kate: figure out how circular this is
- Jim: how much movement do we need while aligning this?
- How well do we align our beam pipe? Doug P: usually try to get it to around 250 microns. But on just a straight beam pipe we don’t usually worry about it so much. On this one, because we’re trying to do the quads, we would really spend the time trying to get it right.
- What Doug’s worried about is if the flanges are off or the vacuum boxes are off
- As long as we have the triplet properly aligned we can compensate somewhat after that.
- Fast to adjust not important; stable is important
- 10% smaller beam pipe? Maybe?
- TODO Kate ask Thomas
- Challenges:
- TODO: test out how well we can align right now, with some shims on a piece of beam pipe. If we actually need more space, and we only need a mm or so, we can look into the smaller pipe
- Doug will talk to Marco
- Where are we at with the model?
- Jim’s working on it
Mapping the magnetic field in the spectrometers
- Needs to be done at TRIUMF
- Ship in 3 weeks ish? Ernie: not sure
- Can we map them in vertical position or do we have to lie them flat?
- Doug P: difficult to get all the way through. Putting them flat is easier.
- What are we measuring? B vs I, or map the entire surface?
- B vs I, MIT can do with a hall probe before they send it
- Then spend the time to do a 3d map of the whole thing. MIT has no capability to do that
- Doug P is a bit concerned about overall length, fringe fields, etc
- Can vacuum chamber go in without splitting the magnets? Should be OK.
- Ernie isn’t sure. Jim isn’t sure
- Jim to look at it tomorrow, and they can test it there. To be answered. Would have to feed in from the top, for sure.
- Tie bar on the magnet comes out, so should be fine.
- Map it when the magnet’s horizontal, and without the vacuum chamber in it.
- Power to magnets while testing:
- 12 weeks from when they ordered the power supplies, which was around last month. So it’ll still be a while. So we need something else we can use.
- Power supplies: interface bus changed but not voltage and current
- Less than 20 feet from power supply to magnet, when testing. So Ernie will tell us what we need. Can run local mode.
- Doug H: should be tested up to equivalent of 25 MeV. Ernie can give us equivalent for current to test at for 10, 15, 20, and 25 MeV.
- Ernie to give Doug P the currents to test at.
- Doug P thinks we can borrow a power supply without too much fuss
- Water: available in both the tents where we might use this. Should be set.
TODO Kate: add these various mappings to the integration checklist
August 13, 2024
Discussion of alignment and PMQs
- Can experiment handle misalignment of beam striking target?
- Doug says lateral shifts should be OK. What about angular shifts? Fraction of a degree.
- If we know where it is, and it’s consistently at that place, we can correct for it.
- Thomas: how consistent?
- Collimator that sets entry into spectrometers is around 400 mm from target. So 1/400 = very small fraction of a degree = 2.5 mRad
- Tells us that lateral offsets will have a bigger impact than angular ones, since 2.5 mRad would be very large for the beam
- Would be good if we know it when taking data because then we can simulate for it, calculate for it, correct for it
- For alignment, let’s see if we can get all 3 PMQs aligned relative to each other on a table in the lab before we take them downstairs and install the whole pipe
- This will also let us see if there’s any cross talk between the quads.
- TODO us: book a week of Shaquille’s time for this, OR convince Marco to let us use the equipment
- This is important ahead of the review, so we can talk about it there.
- For the review
- Steering (Thomas), magnet results with shielding (me)
- Physical layout (Stephi)
- Aveen: show design document
- Stephi chairs
- Aveen’s doc first, then Jim, then alignment/steering/diagnostics etc
- Find PMOG design review form and fill that in (just the usual template I think)
Bellows
- Jim and Joe were looking at the bellows for the target chamber, and it’s basically too tight to work as it is.
- Doug Preddy: they loaded studs into the flange by the bellows, put it in place, then screwed the studs into the threaded thing and nuts in the outside edge. A pain to do, but did allow them to do up the joint
- Joe: even doing that there’s not enough clearance to compress the bellows and get a conflat seal in there.
- [Kate and Thomas went to talk to Petr]
- Easiest solution to the problem is to make a smaller bellows. The ID would be smaller. Say a 1.5 inch bellows on the ID. 2 inch OD and 1.75 inch ID might work.
- Thomas: no.
- Next solution: drill out flange on the dipole?
- Doug Preddy: another idea that I did not understand, but Joe will check
- Jim and Joe will discuss tomorrow and see what they can do
What about the length of the beamline? Has that been fixed now?
- Stephi will follow up with Aveen
- Joe will follow up with Jim
- Doug H has things to follow up with them
Concrete
- TriStar is pretty good and pretty fast these days. Could slip a suitcase block with a hole in it into that order. They have a good form for that already
- Order of magnitude budget estimate: 3k for a 6 foot block
- Drilling holes through the suitcase block could irritate Dragan anyway
July 30, 2024
Motor info for controlling the target ladder
- We have it spoken but we want it written. Please answer email
- Just the little thing that’s on the target ladder. Eastern Air Devices, 34 body stepper motor. Not made anymore. Ernie is checking his filing system
Confirmation of the ion pump purchase - send us the catalog info
- Any other pumps? This is it, right? I got info on power supply and cable from Jim but still missing info on pump itself. Needed for Dimo.
- Ion pump: gamma vacuum Titan ion pump. 2 3/4 conflat. Will dig up recipe for activating it.
Beamline connection - did the 2D drawings work
- TODO Kate: re-send Jim the version of the dump drawing that has the collimator in it
- TODO Kate: schedule review, aiming for 14th & 15th.
Doug has convectron gauge and 175 feet of cable for it in his car. Will bring it out to bates
Adding bellows now, making stands next
July 16, 2024
The model
- Went through the 2D drawings with Jim to figure out what’s what and where it needs to go
Cables
- Power cables from power supplies to spectrometers, from their power supplies. 2 each: positive and negative to the power supply. 244A, unspecified volts. Power supply is 32V to cover DC cable loss, heating, etc. 250 A is DC current from the power supply. Plenty of current and voltage.
- Doug Preddy: we have MCM 300-A cables. We can use those.
- Flow meter and thermal switch cables: one of each for each spectrometer = 4 total. Small easy cables, but they still need to be pulled
- Controls for this: is Jan’s business. So Doug P needs to send the info on this to Jan so he can set it up.
- Also need place (and cords?) to plug them in upstairs?
- Ethernet cables from power supplies to computer/switch (not to downstairs)
- Fiber cable from hut downstairs to switch upstairs. Need
- Ethernet cables from Hall probe controllers to the switch downstairs (in the hut)
- Hall probe cables from controller in the hut to the Hall probes. Probably what it came with is only 2m and we’d need to make new ones, but this is a Doug H question and we need to poke him on that
- Power outlets in the hut [TBD]
- Ethernet cables from VME crate to switch
- Pump power supply
- One being sent by MIT. Has to go on the roof. They have a 15m cable that they can send and we can scavenge the ends off it to use on a longer cable.
- Ernie is going to send info on the other power supply
Need to get five power cables made here at the lab.
- Four for the spectrometer power supplies. Ends on cables? Ernie will tell us.
- Not a long lead time item as long as we have the components. Pull it in then put the ends on.
- Need to know a location for the power supplies before we can know lengths
Magnets
- Done and boxed at Buckley and now Ernie is working on getting them shipped
Screw pitch for the target controller
- 200 steps per revolution
- Linear travel per 1.8 degree step: 5/10000ths of an inch
- Pitch = 0.1 inches per rotation
- 3/8 by 10 ACME thread (book from 1995)
- We should probably measure it anyway, since what Chris has in the model is not representative of that
- Kate has calibration for voltage to position
July 2, 2024
Kate updates
- Problems with electrical group - no progress getting plug design done for hut
- Should get progress on the cables & power supply locations for the spectrometers
- Ion pump power, 2x spectrometer power supplies.
- TRIUMF needs to do spectrometer cables
- Are we doing ion pump power cable? -> Jim to check
Stephi questions & replies
- One water flow connection each for spectrometers -> yes
- What is the vertical distance the ladder moves with one full turn of the screw? -> Jim to ask Chris
- Spectrometer dipole is controlled by experiment? -> yes
- Need to name the spectrometer dipoles so we can tell apart -> call them left and right (looking downstream).
Should look at new PEEK design
Jim will follow up with Joe on whether or not the 2D drawings solved the problem of beam pipe connections between optics & dump
June 18, 2024
Stephi meeting with Mike Rowe from Controls to talk about work request
- Will cover everything we need controls-wise, ladder and whatnot
- Talking also with Doug Preddy about what we need there.
TODO Us: write request document
- To write
- Do not worry if it feasible or not feasible. Say exactly what we need, or want, without safety margins
- Jan: this is difficult because it’s not as simple as that.
- Stephi: great, then write down what we really need and what would be nice. Do a hierarchy. Absolutely necessary, or nice, or these three things together enable this, or whatever. But need to be clear about it.
- Carbon is necessary, tantalum is would be nice.
- Looks like we have to go low energy to high energy
- Only question is do we do something earlier or not.
- If the magnets come very late, and we’re hoping they don’t, then the only reason for running without the magnets in place is to measure background and radiation rates. Put target in and measure the difference. Doug doesn’t really think that’s a necessary measurement, but maybe that’s useful for the machine?
- Thomas: for the machine it’s actually more progress than useful.
Gate reviews
- Thomas reminds us that we need Gate 3 to install, and this means we need to present a solid plan and then do it
Magnets
- Ask Sabr if:
- Wrapper can be flush with the magnets, so as to not increase radius by more than the thickness of the shielding?
- Can the 0.6 magnets get any smaller? Even a mm or two can help?
- Doug thinks that very little mu metal is actually needed. Not 2 mm.
Task for Angela
- Learn how to do field maps for PMQs
- Wrap with mu metal: test what thickness we need so that it works. Thomas will explain what that means. Is 2 mm enough?
- Re-map the magnets equipped with mu metal and see what the effect is on the field integral
- Needs to work for Thomas and Shaquille(?) Thomas will set up a time.
Wait and see what information comes of that first
June 4, 2024
- Cable for ion gauge - already have one so just duplicate it. Cost: $3000, Doug will take a look for a spare
- Alternatively let Jim know what cable we need and he might have it
- BPM upstream -> connect the buttons up
- Pump-out station: something that can start an ion pump. Doug says we have a turbo station on a cart.
- Request to space committee for counting room? Discussion on office space?
Might affect installation: test beam time at DESY at end of October to 11th of November. Ethan going too.
May 21, 2024
To discuss
- Installation schedule
- Approvals
- Commissioning plan
PEEK situation
- At the edge of Jim’s comfort zone, but not bad. Even a couple inches might be OK?
- Optimise for what’s best for the physics and send it to Jim. Takes him less than an hour.
- Laura to send STL files.
Beam pipe
- [TODO] We are going to deliver beam pipe to Bates? How much?
- [TODO] Send Joe collimator specs
- [TODO] Order beam pipe to Bates
- [TODO] Order PMQs to Bates
- TODO for discussion: figure out beam dump end of pipes: reuse something existing, or make pipe all the way to collimator?
- Check with Doug P/Aveen re distances to dump entrance
- [TODO] Get a solidworks file that has the dump in it
Request
- MIT way over budget for DarkLight. Can this be spun off to us?
- TRIUMF is going to buy the long cables
- $15000 USD for power supplies. Can we pay for those?
- Going to be a hard sell.
Installation
-
Need help from TRIUMF
-
Stuff TRIUMF has to do
- Vacate area
- Cabling [where are we getting the tables]
- Concrete hut & tray
-
Magnets
- A month of machining at Bates
- Ship vacuum chambers and magnets together
-
[TODO] Send example commissioning plan to Doug
Will need to swap out what’s on the target ladder: replace thick foil with carbon
10 weeks lead time on IR window
April 17, 2024
Doug: want to outline all the bits and pieces that need to be put in place and who is responsible for which parts. Doug says he will document a lot of this stuff.
Target chamber
-
Stick with the three openings on the ladder
-
Want: 1 micron tantalum and 1 micron carbon. All at TRIUMF. So when assembly is dismounted those need to go in
-
Two windows at beam height looking at the target in position. One is designed to be optical viewing port and the other is thermal. Question to Jim: are the ports on the target chamber we currently have the same size as the ones on the new chamber?
- Flange dimensions on new chamber: 2.75 inch OD. ID: 1.37 inch.
- Flange dimensions on old (currently installed) chamber: identical.
-
Conclusion: Doug would like us to reuse these. We are not convinced that this is the most sensible way to do this, but he seems to be sure it is.
- Ernie and Jim agree with me that it is easier to just buy new windows and install at Bates. Will do.
-
Do we need to take it out? Or can it actually be done in place? Follow up with controls guy.
-
TODO Kate and Stephi: get hold of camera/box
-
Controls: being handled by TRIUMF.
-
Camera: being handled by Stephi
-
Thermal: would be handled by us. And has to go on the port that’s near the wall.
-
We’ll put in the button monitors for the BPM when it arrives.
-
Pipes to the dipoles will be blanked off just upstream of the bellows
-
Doug wants a thin window on the ends of the arms until we get the magnets. Jim will take a look
-
Question about gauges: can we borrow it from TRIUMF or do we have to buy one?
- Buy them, but email Dimo for part numbers first
- Assuming we can order these gauges, or TRIUMF sends a pair, do we need a controller? Jim: yes. TODO Stephi follow up and find out.
-
TODO TRIUMF: figure out cables for gauge controllers as well as magnet power supplies
-
They will make the stand
-
TODO TRIUMF: make the shielding; need design from MIT first
Integrating diagnostics etc into beamline
- Who should Stephi send things to? Joe.
- Joe and Jim will retry opening the step file; it’s very large and slow
- Stephi: considering that when we order the new PMQs
- Ernie: won’t work to wrap in mu metal as a field shield. The mu metal will saturate really quick.
- Confirmed to Ernie that the stray field is problematic
- Ernie: could we add some steel to the yoke, maybe
- Jan: not necessarily linear if you have saturation effects; should we retry at the right field?
- Kate: we actually need to run at 30 sometimes though
- Ernie: if we make the matrix out of something magnetic then Sabr probably won’t be able to get the magnets into it. But ask them.